# EvalNet: A Practical Toolchain for Generation and Analysis of Extreme-Scale Interconnects

### v3

@@ Abstract.
The diversity of communication paths in a network—especially non-minimal paths—is a key enabler of performance at extreme scales. We present EvalNet, a toolchain for scalable generation and analysis over 25 important network topologies, such as Slim Fly, PolarFly, and Orthogonal Fat Trees, with a strong focus on path diversity metrics. EvalNet provides an extensive and fine-grained analysis of shortest and non-shortest paths, including their multiplicities, lengths, and interference. It supports exact measurement and visualization of bandwidth and throughput between every router pair, enabling unprecedented insight into routing potential. EvalNet also includes detailed models for construction cost and power consumption, and interfaces seamlessly with established simulators, which we tune to support large-scale evaluations on low-cost hardware. Using EvalNet, we deliver the widest and most comprehensive path diversity study to date, demonstrating how path diversity underpins throughput and scalability, and facilitating progress towards new frontiers in extreme-scale network design.
copyright: noneccs: Networks Network performance evaluationccs: Networks Network performance analysisccs: Networks Network experimentationccs: Networks Data center networksccs: Networks Packet-switching networks
1. INTRODUCTION
@@
The interconnection network plays an important role in today’s large-scale computing systems (Besta and Hoefler, 2014; Niranjan Mysore et al., 2009; Handley et al., 2017; Valadarsky et al., 2015; Besta et al., 2021; Di Girolamo et al., 2019; Copik et al., 2021; Hoefler et al., 2020; Di Girolamo et al., 2022). Large networks with tens of thousands of nodes are deployed in warehouse-sized HPC and data centers (Fu et al., 2016; Liao et al., 2014). Future supercomputers as well as mega data centers will require even larger scales with hundreds of thousands of servers.
@@
A modern network topology must satisfy many stringent requirements related to different aspects of the interconnect design, such as high performance or low construction cost. However, the evaluation of all these aspects is challenging. First, there are many networks that must be considered for comparison when designing a new topology, for example Fat Tree (Leiserson et al., 1996), Hypercube (Dally and Towles, 2003), Torus (Dally and Towles, 2003), Dragonfly (Kim et al., 2008), Slim Fly (Besta and Hoefler, 2014), Xpander (Ahn et al., 2009), or Jellyfish (Singla et al., 2012) – to name only a few. On top of that, many of these baseline networks, for example Xpander or Orthogonal Fat Trees, do not have straightforward construction algorithms, making it challenging and time-consuming to compare to these baselines.
@@
Similarly, there are many performance measures that should be considered for overall evaluation, and these measures often require different tools. A core challenge in evaluating interconnects is the analysis of path diversity, which has emerged as a critical enabler of performance and scalability in modern networks (Besta et al., 2020). As networks grow to extreme sizes, there is a growing need for fine-grained and scalable tools that can assess not only the existence of diverse paths but also their structure, interference, and utility for multipath routing. In parallel, several complementary performance metrics must also be considered, such as throughput and latency. Throughput is ideally evaluated using packet-level simulators, which offer high accuracy but often do not scale to large topologies. Flow-based simulations scale more easily, but they tend to produce less accurate predictions (Casanova et al., 2008). Furthermore, additional aspects like construction cost and power consumption must be analyzed. Unfortunately, the tools supporting these various metrics often rely on different formats for specifying networks, which significantly complicates integrated and efficient evaluation.
@@
To address all these issues, we propose EvalNet: a publicly available toolchain that facilitates the integrated and comprehensive evaluation of networks. EvalNet consists of several modules, some of which were developed from scratch, while others are interfaces to existing tools which facilitate seamless operation. First, EvalNet enables generation of a very large number of network topologies. We cover the state-of-the-art class of low-diameter topologies developed by optimizing the network structure towards the Moore Bound (Slim Fly (Besta and Hoefler, 2014), PolarFly (Lakhotia et al., 2022), BundleFly (Lei et al., 2020), Xpander (Valadarsky et al., 2015)), Dragonfly topologies (fully-connected Dragonfly (Kim et al., 2008), Cascade Dragonfly (Faanes et al., 2012)), random networks (Jellyfish (Singla et al., 2012)), torus related topologies (Mesh, Torus, Hypercube (Dally and Towles, 2003), Express Mesh (Jain et al., 2017), HyperX (Ahn et al., 2009), Flattened Butterfly (Kim et al., 2007), Tofu (Ajima et al., 2018, 2012)), indirect topologies (Fat Tree (Leiserson et al., 1996), Multi-Layer Full-Mesh (Kathareios et al., 2015), Orthogonal Fat Tree (Kathareios et al., 2015), eXtended Generalized Fat Tree (Ohring et al., 1995), k-ary n-tree (Petrini and Vanneschi, 1997)), and Kautz networks (Undirected Kautz (Li et al., 2004), Arrangement Network (Day and Tripathi, 1992)).
Figure 1. An overview of EvalNet.
[…]
[Figure is fully described in the text.]Figure is fully described in the text.
@@
Second, we observe that the high performance of network workloads is becoming increasingly dependent on effectively harnessing various forms of path diversity, particularly through non-minimal multipathing (Besta et al., 2020). This is especially critical in state-of-the-art low-diameter topologies, such as Slim Fly and PolarFly, which typically offer limited shortest-path diversity but compensate with a rich structure of non-minimal paths. In such networks, the ability to balance load and avoid congestion hinges on the capacity to exploit this non-shortest-path structure. To address this need, EvalNet provides a comprehensive suite of fine-grained tools for analyzing path diversity. It supports the computation of shortest and non-shortest path multiplicities, detailed distributions of path lengths, and interference-aware metrics that capture how routes interact with each other. EvalNet also includes visualization and measurement tools that expose the exact bandwidth and multipathing potential between every pair of routers, enabling both coarse-grained insights and fine-grained optimization of routing and scheduling schemes (Lakhotia et al., 2023). All path diversity analyses are designed to scale to very large topologies, supporting networks with more than 100,000 servers, thus making them practical for evaluating future-generation systems.
@@
Moreover, EvalNet seamlessly interfaces with different tools for network simulation. This includes the flow-level evaluation tool TopoBench (Jyothi et al., 2016) that enables approximating throughput for arbitrary traffic patterns and comes with a specially designed worst-case scenario that puts particularly large stress on the interconnect. EvalNet also harnesses two complementary packet-level simulators (htsim (Raiciu et al., 2010) and Booksim2 (Jiang et al., 2013)), which target, respectively, lossy and lossless architectures [^1] EvalNet does not compete with existing simulators as it does not introduce a new simulator. Instead, it tunes two of these simulators and harnesses them within its full comprehensive pipeline.. We tune these simulators to support large-scale evaluations on low-cost hardware, achieving simulations of more than 1 million endpoint networks on a commodity PC. Finally, EvalNet also comes with a construction cost and power consumption assessment module. The user can parametrize this module with arbitrary cost and power models for both routers and cables, considering arbitrary vendors and architectures such as Ethernet or InfiniBand.
@@
Finally, we use EvalNet to conduct the widest-ever comparative analysis of interconnect topologies, focusing on the path diversity measures, but also combining them with classical metrics that include latency, global bandwidth, construction cost, and power consumption. We investigate path diversity in-depth because other measures (latency, bandwidth under flow/packet simulations, fault-tolerance) have already been extensively researched. However, path diversity has only been analyzed in a limited scope. We provide a total of 55 rich plots for path diversity metrics, combined with observations, analyses, and insights that could be used to develop better routing strategies.
2. EVALNET TOOLCHAIN
@@
We illustrate the overall structure of EvalNet in Figure 1. First, the user selects a topology to be generated; it can be any of the networks described in Section 3; the user can also provide a script generating their own interconnect. The user can then use a module for analyzing the path diversity of a network, which can subsequently be used to enhance routing schemes in simulations (an overview is in Section 4.1, details and evaluation are in Section 5). The generated networks can also be evaluated using either flow or packet-based simulations (overview is in Section 4.2, details and evaluation are in Section 6 and 7, respectively). These simulations are conducted by external tools (TopoBench (Jyothi et al., 2016) for flow-based and htsim (Raiciu et al., 2010) as well as Booksim2 (Jiang et al., 2013) for packet-based simulations of lossy and lossless networks, respectively); we provide input/output conversion routines to integrate these tools seamlessly. Finally, the user can use our cost and power models to assess the construction cost or static power consumption (Section 4.3).
@@
We now proceed to outline these factors (Sections 34) and then detail each of them and provide extensive evaluations (Sections 57). We will finalize with a comprehensive summary comparing all considered interconnects and measures (Section 8).
3. RELEVANT NETWORK TOPOLOGIES
@@
We first describe network topologies included in EvalNet – see also Figure 2. The most important symbols are listed in Table 1.
@@
$N,N_{r}$endpoints and #routers in the network ($N_{r}=|V|$).
$p,k^{\prime}$endpoints attached to a router, #channels to other routers.
@@
Table 1. The most important symbols used in this work.
Figure 2. An overview of network topologies offered by default in EvalNet. New topologies can be straightforwardly added to the pipeline.
[…]
[Figure is fully described in the text.]Figure is fully described in the text.
3.1. Network Model
@@
While EvalNet supports arbitrary configurations of $p$, $k$, and $k^{\prime}$, in the experiments, we usually set $p=\frac{k^{\prime}}{d}$, as it maximizes throughput while minimizing congestion and network cost (assuming a random uniform traffic pattern) (Besta et al., 2020). We also usually select network radix $k^{\prime}$ and router count $N_{r}$ so that, for a fixed $N$, the compared topologies use similar amounts of networking hardware and thus have similar construction costs. We assume router-based topologies hereafter, i.e., no direct channels between any two endpoints.
3.2. Moore Bound Topologies
A recent class of low-diameter topologies explicitly optimizes the network structure towards the Moore Bound (Besta and Hoefler, 2014). This maximizes the number of routers and endpoints for a fixed diameter and radix, minimizing cost and power consumption. These topologies are important as they have achieved new frontiers in performance, cost, and power consumption.
Slim Fly (SF) (Besta and Hoefler, 2014), PolarFly (PF) (Lakhotia et al., 2022), and PolarStar (Lakhotia et al., 2024) belong to this class, they are based on the underlying McKay-Miller-Širáň and Erdős-Rényi polar graphs, respectively. Spectralfly (SpF) (Young et al., 2022) is based on the Ramanujan graph construction of Lubotzky, Phillips, and Sarnak (LPS). Moreover, a $\tau$-regular Xpander (XP) (Valadarsky et al., 2015) constructions starts with a complete graph of $\tau+1$ vertices, and is then repeatedly “lifted” ($\frac{N_{r}}{\tau+1}$ times) to obtain the final graph with good expansion properties. Finally, BundleFly (BF) (Lei et al., 2020) is using SF together with a Paley graph. The routers in both graphs are linked via a multi-star product based on a bijection.
3.3. Dragonfly Topologies
The Fully-Connected Dragonfly (DF) (Kim et al., 2008) is comprised of $g$ groups of nodes internally interconnected in an all-to-all fashion. The topology among groups is an all-to-all, too. A Cascade Dragonfly (C-DF) (Faanes et al., 2012) alters the original Dragonfly by assuming a 16x6 2D HyperX for intra-group and an all-to-all graph for inter-group connectivity. Megafly (MF) (Flajslik et al., 2018) is another Dragonfly variant. These networks are relevant as they come with good performance and cost, albeit achieving worse values for these metrics than low-diameter networks, but have been found to be easier to cable. They also offer relatively easy scheduling: their modular structure facilitates achieving data locality for workloads (Tate et al., 2014).
3.4. Random Topologies
Fully random networks are infeasible due to the unbounded radix. The Jellyfish (Singla et al., 2012) is a random graph with a pre-selected router radix $k$. We support these networks as they come with advantageous performance, cost, and power consumption (Singla et al., 2012; Besta et al., 2020).
@@ 3.4. Random Topologies
Figure 3. An overview of evaluation measures of network topologies offered by default in EvalNet. New measures can be straightforwardly added to the pipeline.
[…]
[Overview of evaluation measures for network topologies.]The figure provides an overview of the structure of the Evaluating Network Topologies chapter as well as the considered metrics. The chapter consists of path diversity measures (4.1), cost and power (4.3) as well as throughput and latency (4.2) subsections. The last subsection considers throughput (saturation for a given workload), maximum achievable throughput (MAT), network latency and average path length. The second subsection looks at the construction cost and power consumption. The path diversity subsection investigates counts for minimal and non-minimal paths as well as path interference and total network load (TNL). For the minimal and non-minial paths specifically lengths and multiplicity are considered.
3.5. Torus Related Topologies
@@
We consider traditional Mesh (MH), Torus (TR), and Hypercube (HC) topologies. We also incorporate their more recent enhancements: HyperX (HX) (Ahn et al., 2009) (each router in $n$-D grid connects to all other routers within the same dimension), Express Meshes (EM) (Jain et al., 2017) (cut-down versions of the HyperX), Flattened Butterfly (FBF) (Kim et al., 2007) (a 2D variation of HyperX), and Tofu (TF) (Ajima et al., 2018, 2012) (a 6D torus variant with inner 3D mesh/tori networks embedded as virtual routers). These networks have been widely deployed across different settings and thus are considered in EvalNet.
3.6. Indirect Topologies
@@
In indirect networks, some routers attach endpoints while other routers only forward packets. A Fat Tree (FT) (Leiserson et al., 1996) is a binary tree, with either higher bandwidth links towards the root or multiple parallel links towards the root to avoid communication bottlenecks. It is one of the most widely researched and used network topologies (Katta et al., 2016b; Vanini et al., 2017; Alizadeh et al., 2014; Katta et al., 2016a; He et al., 2015; Zhang et al., 2017; Huang et al., 2020; Alasmar et al., 2018; Montazeri et al., 2018; Hwang et al., 2014; Lu, 2016; Vamanan et al., 2012; Alizadeh et al., 2013; Bai et al., 2014; Raiciu et al., 2011; Handley et al., 2017; Zhuo et al., 2016; He et al., 2016; Alizadeh et al., 2010; Cardwell et al., 2017; Mittal et al., 2015; Li and Pan, 2013; Perry et al., 2014; Olteanu and Raiciu, 2016; Park et al., 2019; Benet et al., 2018; Lu et al., 2018; Widmer et al., 2001; Callegari et al., 2014). The k-ary n-tree (AT) (Petrini and Vanneschi, 1997) avoids infeasible high-radix roots of FT by using the same number of routers per tree-level and connecting them with increasing strides towards the top. An eXtended Generalized Fat Tree (xGFT) (Ohring et al., 1995) is a k-ary n-tree variant which allows reducing deployment cost. Orthogonal Fat-Trees (OFT) (Kathareios et al., 2015) are based on k-ary 2-trees; the wiring pattern between the two levels is modified to increase scalability. Finally, a Multi-Layer Full-Mesh (MLFM) (Kathareios et al., 2015) consists of multiple 2D meshes (incl. diagonal connections) stacked together, but instead of direct links it uses global routers to link meshes. EvalNet also supports indirect networks as they offer very high bandwidth; FTs are also widely deployed.
3.7. Kautz Topologies

### v4

@@ Abstract
The diversity of communication paths in a network—especially non-minimal paths—is a key enabler of performance at extreme scales. We present EvalNet, a toolchain for scalable generation and analysis of over 25 important network topologies, such as Slim Fly, PolarFly, and Orthogonal Fat Trees, with a strong focus on path diversity metrics. EvalNet provides an extensive and fine-grained analysis of shortest and non-shortest paths, including their multiplicities, lengths, and interference. It supports exact measurement and visualization of bandwidth and throughput between every router pair, enabling unprecedented insight into routing potential. EvalNet also includes detailed models for construction cost and power consumption, and interfaces seamlessly with established simulators, which we tune to support large-scale evaluations on low-cost hardware. Using EvalNet, we deliver the widest and most comprehensive path diversity study to date, demonstrating how path diversity underpins throughput and scalability, and facilitating progress towards new frontiers in extreme-scale network design.
I INTRODUCTION
@@
The interconnection network plays an important role in today’s large-scale computing systems [15, 78, 42, 100, 13, 32, 28, 45, 31]. Large networks with tens of thousands of nodes are deployed in warehouse-sized HPC and data centers [37, 72]. Future supercomputers as well as mega data centers will require even larger scales with hundreds of thousands of servers.
@@
A modern network topology must satisfy many stringent requirements related to different aspects of the interconnect design, such as high performance or low construction cost. However, the evaluation of all these aspects is challenging. First, there are many networks that must be considered for comparison when designing a new topology, for example Fat Tree [69], Hypercube [29], Torus [29], Dragonfly [63], Slim Fly [15], Xpander [1], or Jellyfish [94] – to name only a few. On top of that, many of these baseline networks, for example Xpander or Orthogonal Fat Trees, do not have straightforward construction algorithms, making it challenging and time-consuming to compare to these baselines.
@@
Similarly, there are many performance measures that should be considered for overall evaluation, and these measures often require different tools. A core challenge in evaluating interconnects is the analysis of path diversity, which has emerged as a critical enabler of performance and scalability in modern networks [17]. As networks grow to extreme sizes, there is a growing need for fine-grained and scalable tools that can assess not only the existence of diverse paths but also their structure, interference, and utility for multipath routing. In parallel, several complementary performance metrics must also be considered, such as throughput and latency. Throughput is ideally evaluated using packet-level simulators, which offer high accuracy but often do not scale to large topologies. Flow-based simulations scale more easily, but they tend to produce less accurate predictions [23]. Furthermore, additional aspects like construction cost and power consumption must be analyzed. Unfortunately, the tools supporting these various metrics often rely on different formats for specifying networks, which significantly complicates integrated and efficient evaluation.
@@
To address all these issues, we propose EvalNet: a publicly available toolchain that facilitates the integrated and comprehensive evaluation of networks (contribution 1). EvalNet consists of several modules, some of which were developed from scratch, while others are interfaces to existing tools which facilitate seamless operation. First, EvalNet enables generation of a very large number of network topologies. We cover the state-of-the-art class of low-diameter topologies developed by optimizing the network structure towards the Moore Bound (Slim Fly [15], PolarFly [65], BundleFly [68], Xpander [100]), Dragonfly topologies (fully-connected [63], Cascade [34]), random networks (Jellyfish [94]), torus-related topologies (Mesh, Torus, Hypercube [29], Express Mesh [54], HyperX [1], Flattened Butterfly [64], Tofu [3, 4]), indirect topologies (Fat Tree [69], Multi-Layer Full-Mesh [60], Orthogonal Fat Tree [60], eXtended Generalized Fat Tree [79], k-ary n-tree [84]), and Kautz networks (Undirected Kautz [70], Arrangement Network [30]).
Figure 1: An overview of EvalNet.
@@
Second, we observe that the high performance of network workloads is becoming increasingly dependent on effectively harnessing various forms of path diversity, particularly through non-minimal multipathing [17, 65]. This is especially critical in state-of-the-art low-diameter topologies, such as Slim Fly and PolarFly, which typically offer limited shortest-path diversity but compensate with a rich structure of non-minimal paths. In such networks, the ability to balance load and avoid congestion hinges on the capacity to exploit this non-shortest-path structure. To address this need, EvalNet provides a comprehensive suite of fine-grained tools for analyzing path diversity (contribution 2). It supports the computation of shortest and non-shortest path multiplicities, detailed distributions of path lengths, and interference-aware metrics that capture how routes interact with each other. EvalNet also includes visualization and measurement tools that expose the exact bandwidth and multipathing potential between every pair of routers, enabling both coarse-grained insights and fine-grained optimization of routing and scheduling schemes [66]. All path diversity analyses are designed to scale to very large topologies, supporting networks with more than 1 million servers, thus making them practical for evaluating future systems.
@@
We design EvalNet to enable simulations of massive networks with different architectures even on low-cost commodity hardware (contribution 3). For this, EvalNet harnesses two complementary packet-level simulators (htsim [91] and Booksim2 [56]), which target, respectively, lossy and lossless architectures (EvalNet does not compete with existing simulators as it does not introduce a new simulator. Instead, it tunes two of these simulators and harnesses them within its full comprehensive pipeline). We tune these simulators to support large-scale evaluations on low-cost hardware, achieving simulations of networks with more than 1 million endpoints on a commodity PC with just 16GB of RAM without impacting the simulation accuracy. On top of that, EvalNet also seamlessly interfaces with the flow-level evaluation tool TopoBench [57] that enables approximating throughput for arbitrary traffic patterns and comes with a specially designed worst-case scenario that puts particularly large stress on the interconnect. Finally, EvalNet also provides a construction cost and power consumption assessment module. The user can parametrize this module with arbitrary cost and power models for both routers and cables, considering arbitrary vendors and architectures such as Ethernet or InfiniBand.
@@
Finally, we use EvalNet to conduct the widest-ever comparative analysis of interconnect topologies (contribution 4), focusing on the path diversity measures, but also combining them with classical metrics that include latency, global bandwidth, construction cost, and power consumption. We investigate path diversity in-depth because other measures (latency, bandwidth under flow/packet simulations, fault-tolerance) have already been extensively researched. However, path diversity has only been analyzed in a limited scope. We provide a total of 55 plots rich in path diversity information based on 4 different metrics, combined with observations, analyses, and insights that can be used to develop better routing strategies.
Figure 2: An overview of network topologies offered by default in EvalNet. New topologies can be straightforwardly added to the pipeline.
II EVALNET TOOLCHAIN
@@
Figure 1 illustrates the overall structure of EvalNet. First, the user selects a topology to be generated; it can be any of the networks described in Section III; the user can also provide a script generating their own interconnect. The user can then use a module for analyzing the path diversity of a network, which can subsequently be used to enhance routing schemes in simulations (an overview is in Section IV-A, details and evaluation are in Section V). The generated networks can also be evaluated using either flow or packet-based simulations (overview is in Section IV-B, details and evaluation are in Section VI and VII, respectively). These simulations are conducted by external tools (TopoBench [57] for flow-based and htsim [91] as well as Booksim2 [56] for packet-based simulations of lossy and lossless networks, respectively); we provide input/output conversion routines to integrate these tools seamlessly. Finally, the user can use our cost and power models to assess the construction cost or static power consumption (Section IV-C).
@@
We now proceed to outline these factors (Sections IIIIV) and then detail each of them and provide extensive evaluations (Sections VVII). We will conclude with a comprehensive summary comparing all considered interconnects and measures (Section VIII).
III RELEVANT NETWORK TOPOLOGIES
@@
We first describe network topologies included in EvalNet – see also Figure 2. The most important symbols are listed in Table I.
TABLE I: The most important symbols used in this work.
@@
$N,N_{r}$ #endpoints and #routers in the network ($N_{r}=|V|$).
$p,k^{\prime}$ #endpoints attached to a router, #channels to other routers.
@@
III-A Network Model
@@
While EvalNet supports arbitrary configurations of $p$, $k$, and $k^{\prime}$, in the experiments, we usually set $p=\frac{k^{\prime}}{d}$, as it maximizes throughput while minimizing congestion and network cost (assuming a random uniform traffic pattern) [17]. We also usually select network radix $k^{\prime}$ and router count $N_{r}$ so that, for a fixed $N$, the compared topologies use similar amounts of networking hardware and thus have similar construction costs. We assume router-based topologies hereafter, i.e., no direct channels between any two endpoints.
@@
Figure 3: An overview of evaluation measures of network topologies offered by default in EvalNet. New measures can be straightforwardly added to the pipeline.
III-B Moore Bound Topologies
A recent class of low-diameter topologies explicitly optimizes the network structure towards the Moore Bound [15]. This maximizes the number of routers and endpoints for a fixed diameter and radix, minimizing cost and power consumption. These topologies are important as they have achieved new frontiers in performance, cost, and power consumption.
Slim Fly (SF) [15], PolarFly (PF) [65], and PolarStar [67] belong to this class, they are based on the underlying McKay-Miller-Širáň and Erdős-Rényi polar graphs, respectively. Spectralfly (SpF) [106] is based on the Ramanujan graph construction of Lubotzky, Phillips, and Sarnak (LPS). Moreover, a $\tau$-regular Xpander (XP) [100] construction starts with a complete graph of $\tau+1$ vertices, and is then repeatedly “lifted” ($\frac{N_{r}}{\tau+1}$ times) to obtain the final graph with good expansion properties. Finally, BundleFly (BF) [68] is using SF together with a Paley graph. The routers in both graphs are linked via a multi-star product based on a bijection.
III-C Dragonfly Topologies
The Fully-Connected Dragonfly (DF) [63] is comprised of $g$ groups of nodes internally interconnected in an all-to-all fashion. The topology among groups is an all-to-all, too. A Cascade Dragonfly (C-DF) [34] alters the original Dragonfly by assuming a 16x6 2D HyperX for intra-group and an all-to-all graph for inter-group connectivity. Megafly (MF) [36] is another Dragonfly variant. These networks are relevant as they come with good performance and cost, albeit achieving worse values for these metrics than low-diameter networks, but have been found to be easier to cable. They also offer relatively easy scheduling: their modular structure facilitates achieving data locality for workloads [98].
III-D Random Topologies
Fully random networks are infeasible due to the unbounded radix. The Jellyfish [94] is a random graph with a pre-selected router radix $k$. We support these networks as they come with advantageous performance, cost, and power consumption [94, 17].
III-E Torus-Related Topologies
@@
We consider traditional Mesh (MH), Torus (TR), and Hypercube (HC) topologies. We also incorporate their more recent enhancements: HyperX (HX) [1] (each router in $n$-D grid connects to all other routers within the same dimension), Express Meshes (EM) [54] (cut-down versions of the HyperX), Flattened Butterfly (FBF) [64] (a 2D variation of HyperX), and Tofu (TF) [3, 4] (a 6D torus variant with inner 3D mesh/tori networks embedded as virtual routers). These networks have been widely deployed across different settings and thus are considered in EvalNet.
III-F Indirect Topologies
@@
In indirect networks, some routers attach endpoints while other routers only forward packets. A Fat Tree (FT) [69] is a binary tree, with either higher bandwidth links towards the root or multiple parallel links towards the root to avoid communication bottlenecks. It is one of the most widely researched and used network topologies [62, 102, 6, 61, 43, 109, 48, 5, 76, 49, 73, 101, 8, 9, 90, 42, 111, 44, 7, 22, 75, 71, 83, 80, 81, 11, 74, 104, 21]. The k-ary n-tree (AT) [84] avoids infeasible high-radix roots of FT by using the same number of routers per tree-level and connecting them with increasing strides towards the top. An eXtended Generalized Fat Tree (xGFT) [79] is a k-ary n-tree variant which allows reducing deployment cost. Orthogonal Fat-Trees (OFT) [60] are based on k-ary 2-trees; the wiring pattern between the two levels is modified to increase scalability. Finally, a Multi-Layer Full-Mesh (MLFM) [60] consists of multiple 2D meshes (incl. diagonal connections) stacked together, but instead of direct links it uses global routers to link meshes. EvalNet also supports indirect networks as they offer very high bandwidth; FTs are also widely deployed.
III-G Kautz Topologies